OpenStudy (anonymous):

solve for n P(n,3) = 60

OpenStudy (luigi0210):

I'm assuming the P means Permutations.. do you know the formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

P(n,r) = n!/(n-r)!

OpenStudy (luigi0210):

So I'm assuming they want us to work backwards.. \(\large P(n, r ) = 60 = \frac{n!}{(n-3)!}\) And we know that that \(n> 3 \)

OpenStudy (anonymous):

yeah, for sure :)

OpenStudy (luigi0210):

I don't think they'd make the problem too difficult.. so I'd suggest just plugging in values for n that are greater than 3. So if we try 4; 4! would equal 24, which is too small and not equal to 60. So 4 won't work. Now trying 5: \(\large P(5, 3) = \frac{5!}{(5-3)!} = \frac{120}{2!}\) And what's 120/2?

OpenStudy (anonymous):

how to solve this then? i'm required to show a solution

OpenStudy (anonymous):

\[0=n ^{3}-3n ^{2}+2n-60\]

OpenStudy (anonymous):

i have went all until this equation, and i'm stuck

OpenStudy (anonymous):

but 5's gonna work for that equation for sure

OpenStudy (luigi0210):

I'm not sure if they want you to go into it that deep, I think just plugging in values would suffice

OpenStudy (anonymous):

yeah they do! and it's HORRIBLE

OpenStudy (anonymous):

but thanks anyway :)